bryan bought a new tv for $725 the sales tax is 7% what is the total he paid
Answer by mananth(16946)


bryan bought a new tv for $725 the sales tax is 7% what is the total he paid
sales tax is 7% =7/100 = 0.07
Toatal paid = 725*1.07
- $ 775.75